IF ALEXIS HAS $345 IN HER SAVINGS ACCOUNT AND WILL DEPOSIT $15 EACH WEEK.AFTER HOW MANY WEEKS WILL SHE HAVE $1,035?
Answer by checkley75(3666)

345+15X=1035
15X=1035-345
15X=690
X=690/15
X=46 MONTHS IS REQUIRED TO INCREASE THE 345 TO 1035.